OpenStudy (anonymous):

How can x2 = 2x + 4 be set up as a system of equations?

OpenStudy (anonymous):

I really appreciate you're help!!!

OpenStudy (anonymous):

y = x2 − 2x + 4 y = x2 + 2x + 4 y = x2 y = x2 + 2x + 4 y = x2 + 2x y = x2 + 4 y = x2 y = 2x + 4

OpenStudy (triciaal):

not sure of the system of eqn but to factor rewrite to get x^2 -2x-4 =0 x

OpenStudy (anonymous):

thanks for the help girl!